# Leave-One-Out Neighborhood Smoothing for Graphons: Berry-Esseen Bounds, Confidence Intervals, and Honest Tuning

### 28 Feb 2026 v4

@@ Abstract
Estimating latent edge probabilities in network models often relies on aggregating information from structurally similar nodes. However, using the same adjacency matrix to both determine node similarities and estimate probabilities induces complex statistical dependencies. This coupling invalidates standard inferential tools, such as concentration inequalities and central limit theorems, leaving a critical gap in entrywise uncertainty quantification for networks. To resolve this, we introduce a leave-one-out (LOO) decoupling strategy. When predicting a target edge, the local neighborhood is constructed from a reduced adjacency matrix with the target row and column entirely removed. This structural modification-rather than the specific smoothing mechanism-is the core innovation: it restores exact conditional independence between the data-driven neighborhood selection and the observed edges being aggregated. Leveraging this restored independence, we derive variance-adaptive empirical Bernstein bounds and establish uniform Berry-Esseen normal approximations with explicit rates. These results enable finite-sample, variance-adaptive confidence intervals for the localized neighborhood mean and entrywise uncertainty quantification for $P_{ij}$ on a high-probability interior (good) set up to an explicit bias term. We also establish a uniform Berry–Esseen approximation for the variance component (normal‑approximation intervals use a heuristic variance plug‑in in this version).
Furthermore, the LOO architecture naturally admits an honest cross-validation scheme for tuning parameter selection, providing a complete and rigorous inferential framework while preserving strong mean-squared error performance.
@@ 1 Introduction
Graphon models have emerged as the canonical nonparametric framework for analyzing complex network data, providing a flexible representation of latent edge probabilities through a measurable kernel on the unit square (Borgs et al., 2008; Lovász, 2012). Over the past decade, a rich literature has developed around estimating the underlying probability matrix $P$ or the graphon $f$ from a single observed adjacency matrix $A$. Early foundational work focused heavily on the Stochastic Block Model (SBM) and its variants, utilizing modularity maximization, maximum likelihood, and spectral clustering to recover block probabilities (Bickel and Chen, 2009; Abbe, 2017; Rohe et al., 2011). As the field moved toward continuous, nonparametric graphons, modern estimators-including universal singular value decomposition (Chatterjee, 2015), sorting-based smoothing (Chan and Airoldi, 2014), and block-approximations (Gao et al., 2015)-have been proven to reliably achieve minimax-optimal rates for point estimation.
Among these, the neighborhood smoothing method introduced by Zhang et al. (2017) stands out. By adaptively aggregating structurally similar nodes based on empirical two-hop distances, neighborhood smoothing achieves optimal mean-squared error (MSE) without requiring the strict structural assumptions of block models or the selection of a global low-rank embedding.
@@
This decoupling is the central theoretical lever of the paper. By strictly separating the data used for neighborhood selection from the data used for estimation, we restore exact conditional independence among the centered summands. This topological decoupling permits the application of classical empirical process tools to graphon neighborhood smoothing. Specifically, under standard piecewise Lipschitz assumptions on the graphon, we leverage this restored independence to establish three main contributions:
@@
    Asymptotic Normality: We establish a uniform Berry–Esseen bound with explicit rates for the normalized variance component; normal‑approximation intervals in this version use a heuristic variance estimate and incur an additional variance‑estimation error term.
@@
    Honest Cross-Validation: The LOO structure naturally supports an unbiased cross-validation scheme for tuning the neighborhood size $h_{n}$.inequality. Here, honest cross-validation means that it does not reuse held-out edges in predictor construction, ensuring valid inference even after data-driven tuning.
@@
The remainder of the paper is organized as follows. Section 2 formalizes the graphon model, the Zhang-Levina-Zhu (ZLZ), Zhang et al. (2017) empirical distance, and our leave-one-out neighborhood construction. Section 3.2 establishes the foundational conditional independence, proving the variance-adaptive concentration bounds, Berry-Esseen normal approximations, and the validity of the resulting confidence intervals. Section 4 details the honest cross-validation scheme for tuning parameter selection. Section 5 reports comprehensive simulation results illustrating the finite-sample coverage and MSE performance, and Section 6 concludes with a discussion of extensions to directed, weighted, and degree-corrected settings.
@@ 2 Model and Leave-One-Out Neighborhood Smoothing
We observe a single undirected simple graph on $n$ vertices with adjacency matrix $A\in\{0,1\}^{n\times n}$, no self‑loops ($A_{ii}=0$), and independent upper‑triangular entries
@@
Assumption 2.2 places the model in the dense graphon regime, where edge probabilities are uniformly bounded away from zero and one. This excludes sparse scalings of the form $P_{ij}=\rho_{n}W(\xi_{i},\xi_{j})$ with $\rho_{n}\to 0$. Our Berry-Esseen and empirical Bernstein arguments rely critically on variance bounded away from zero. Extending the present theory to sparse or degree-corrected regimes would require variance renormalization and additional control of small-probability events, and is left for future work.
@@
Assumption 2.2 requires that there exists $\eta\in(0,1/2)$ such that $\eta\leq f(u,v)\leq 1-\eta$ for all $(u,v)\in[0,1]^{2}$. This nondegeneracy condition guarantees a uniformly nonvanishing conditional variance and is used to apply Berry-Esseen and variance‑adaptive concentration bounds. It excludes extremely sparse regimes where $P_{ij}\to 0$ uniformly; extending our results to such regimes requires variance re‑normalization and additional control of rare events, which we leave to future work. Empirically, moderate thinning of edges does not invalidate the procedure (see Section 5).
@@
For concreteness we work in the canonical (dense) graphon regime $P_{ij}=f(\xi_{i},\xi_{j})$ with $f$ fixed as in Assumption 2.1. In this regime the expected degree satisfies $\mathbb{E}\deg(i)=\sum_{j=1}^{n}P_{ij}\asymp n$, so local neighborhoods contain many informative two‑hop paths and several concentration steps simplify. All main results below are stated under Assumptions 2.1 and 2.2 together with this dense regime convention. Extensions to explicit sparsity scalings $P_{ij}=\rho_{n}f(\xi_{i},\xi_{j})$ are discussed briefly in Section 6.
@@ 2.1 LOO neighborhoods and estimator
Define the leave-one-out neighborhood of node $i$ by
$\mathcal{N}_{i}^{(-j)}=\left\{k\neq i,j:d_{ik}^{(-j)}\leq d_{i,(h_{n})}^{(-j)}\right\},$ (2.3)
where $h_{n-1}$ is the ZLZ bandwidth parameter with $n$ replaced by n1n-1. that is, the $h_{n}$ nearest neighbors of $i$ according to the ZLZ two-hop distance computed on the reduced graph.
Given $\mathcal{N}_{i}^{(-j)}$, define the one‑sided LOO smoother and its symmetrized version by
@@
This construction preserves the original ZLZ neighborhood rule, with the sole modification that node $j$ is removed prior to distance computation.
The classical neighborhood smoother of Zhang et al. (2017) uses $\mathcal{N}_{i}$ computed from the full adjacency matrix $A$ to estimate the entire $i$th row of $P$. Define the distance function $d:V\times V\to\mathbb{R}_{+}$ by
@@
This is the sup‑norm discrepancy between the $i$th and $j$th rows of the normalized two‑hop matrix $A^{2}/n$, measuring the similarity of their two‑hop neighborhoods. Therefore the ZLZ neighborhood $\mathcal{N}_{i}$ consists of the $h_{n}$ nodes with the smallest ZLZ distance to $i$. Zhang et al. (2017) defined
$\mathcal{N}_{i}=\{k\in V\setminus\{i\}:d_{ZLZ}(i,k)\leq q_{h}(\{D\}),\quad D=\{d_{ZLZ}(i,k):k\in V\setminus\{i\}\}\},$ (2.6)
and $q_{h}(\{D\})$ is the $h_{n}$-th order statistic of the set $\{d_{ZLZ}(i,k):k\in V\setminus\{i\}\}$. The ZLZ estimator is then defined by
$\widetilde{P}^{\mathrm{ZLZ}}_{ij}:=\frac{1}{|\mathcal{N}_{i}|}\sum_{k\in\mathcal{N}_{i}}A_{kj}.$
Because $\mathcal{N}_{i}$ depends on all columns of $A$, including column $j$, the weights used to average $\{A_{kj}:k\in\mathcal{N}_{i}\}$ depend on the very summands being averaged. Hence the centered terms $A_{kj}-P_{kj}$ are not independent conditional on the neighborhoods, preventing the direct application of variance‑adaptive concentration or Berry-Esseen theory.
@@
The symmetrized estimator $\widehat{P}_{ij}$ respects the symmetry of the simple graph model. All concentration and normal‑approximation results are derived for the one‑sided quantities $\tilde{P}_{ij}$; symmetrization does not affect the independence structure.
@@
Condition on the latent positions $(\xi_{1},\dots,\xi_{n})$ so that $P$ is deterministic with $P_{kl}=f(\xi_{k},\xi_{\ell})$. Under the model (2.1), the edges $\{A_{k\ell}:1\leq k<\ell\leq n\}$ are independent Bernoulli random variables with parameters $P_{k\ell}$.
By construction, $A^{(-j)}$ is obtained from $A$ by removing the $j$th row and column. Hence $A^{(-j)}$ is a measurable function of $\{A_{k\ell}:k\neq j,\ell\neq j\}$, and so is $\mathcal{N}_{i}^{(-j)}$. Therefore, conditional on $P$, the random set $\mathcal{N}_{i}^{(-j)}$ is independent of the collection $\{A_{kj}:k\neq j\}$.
@@
For each $k\in\mathcal{N}_{i}^{(-j)}$, we have $A_{kj}\sim\mathrm{Bernoulli}(P_{kj})$ and $A_{kj}-P_{kj}\in[-1,1]$ with mean zero. Since the edges $\{A_{kj}:k\neq j\}$ are independent conditional on $P$, and $\mathcal{N}_{i}^{(-j)}$ is independent of these edges, it follows that the collection $\{A_{kj}-P_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ is independent, mean-zero, and bounded in [1,1][-1,1] conditional on $(\mathcal{N}_{i}^{(-j)},P)$. ∎
3 Consistency
Our objective is entrywise inference for the finite-sample edge probability matrix $P=(P_{ij})_{1\leq i,j\leq n}$, where $P_{ij}=W(\xi_{i},\xi_{j})$. All probability statements in Sections 3 and 4 are conditional on the realized latent positions $(\xi_{1},\dots,\xi_{n})$, equivalently on $P$. Thus our results provide conditional finite-sample and asymptotic inference for $P_{ij}$ rather than for the graphon $W$ itself. In this section we quantify the deviation of the LOO estimator $\tilde{P}_{ij}$ from the target $P_{ij}$. The analysis separates the estimation error into a variance component, arising from the randomness of observed edges given $P$, and a bias component, arising from approximating $P_{ij}$ by a local average over a neighborhood of $i$. We begin with a decomposition that will be used throughout the section. For $i\neq j$ write For $i\neq j$, write
$\Delta_{ij}:=\tilde{P}_{ij}-P_{ij}=U_{ij}^{(-j)}+B_{ij}^{(-j)},$ (3.1)
where
$U_{ij}^{(-j)}$ $:=\tilde{P}_{ij}-\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(A_{kj}-P_{kj}),$ (3.2)
By Proposition 2.7, conditional on $\mathcal{N}_{i}^{(-j)}$ and $P$ the summands in $U_{ij}^{(-j)}$ are independent, mean zero and bounded, so $U_{ij}^{(-j)}$ behaves like an average of $h_{n}$ independent observations. The term $B_{ij}^{(-j)}$ is a deterministic (given $P$ and $\mathcal{N}_{i}^{(-j)}$) smoothness bias that can be controlled using Assumption 2.1 and the neighborhood properties in Section 2. The following two subsections analyze these two components separately.
3.1 Neighborhood bias
The bias $B_{ij}$ captures the approximation error from replacing $P_{ij}$ by the local average of $P_{kj}$ over the LOO neighborhood. Under the piecewise Lipschitz graphon assumption and the ZLZ neighborhood construction, neighbors in $\mathcal{N}_{i}^{(-j)}$ are close in latent position to $\xi_{i}$, which yields a small bias.
Lemma 3.1 (Uniform two-hop stability under deletion of one vertex).
Let $A\in\{0,1\}^{n\times n}$ be the adjacency matrix of an undirected simple graph with $A_{ii}=0$. For each $j\in[n]$, let $A^{(-j)}$ be the $(n-1)\times(n-1)$ adjacency matrix obtained by deleting the $j$th row and column. Define the normalized two-hop matrices
$M:=\frac{1}{n}A^{2},\qquad M^{(-j)}:=\frac{1}{n-1}\big(A^{(-j)}\big)^{2}.$
To compare $M$ and $M^{(-j)}$, embed $M^{(-j)}$ into an $n\times n$ matrix $M^{(-j)}_{\mathrm{embed}}$ by
$M^{(-j)}_{\mathrm{embed}}(i,\ell)=\begin{cases}M^{(-j)}_{i\ell},&i,\ell\neq j,\\ M_{i\ell},&\text{if }i=j\text{ or }\ell=j.\end{cases}$
Thus $M^{(-j)}_{\mathrm{embed}}$ and $M$ have the same dimensions and agree on the $j$th row and column.
Then, deterministically,
$\sup_{1\leq j\leq n}\ \|M-M^{(-j)}_{\mathrm{embed}}\|_{\max}\;\leq\;\frac{2}{n},$
$|M_{i\ell}-M^{(-j)}_{i\ell}|\leq\frac{2}{n}.$

### 14 Mar 2026 v5

@@ Abstract
Estimating latent edge probabilities from a single observed network is a fundamental challenge in statistical network analysis. While classical neighborhood smoothing methods provide consistent point estimates, providing valid entrywise uncertainty quantification remains a complex problem due to the intricate dependencies formed during data-driven neighborhood selection. In this paper, we propose a novel Leave-One-Out (LOO) neighborhood smoothing estimator that explicitly decouples these dependencies. By strictly excluding the target node’s column during neighborhood formation, our approach unlocks exact conditional independence while preserving the statistical efficiency of the local neighborhood. We rigorously establish the bias-variance trade-off of this estimator by deriving a conditional risk bound, demonstrating that the stochastic cross-term vanishes identically. Leveraging this geometric conditional independence, we construct valid entrywise confidence intervals through two distinct frameworks: strictly non-asymptotic empirical Bernstein bounds for finite-sample guarantees, and a Central Limit Theorem for sharp asymptotic inference under an undersmoothing regime. Extensive simulation studies across diverse graphon topologies confirm that our asymmetric LOO estimator matches the point estimation performance of classical methods while successfully achieving nominal coverage rates for entrywise uncertainty quantification.
@@ 1 Introduction
Graphon models have emerged as the canonical nonparametric framework for analyzing complex network data, providing a flexible representation of latent edge probabilities through a measurable kernel on the unit square (Borgs et al., 2008; Lovász, 2012). Over the past decade, a rich literature has developed around estimating the underlying probability matrix $P$ or the graphon $f$ from a single observed adjacency matrix $A$. Early foundational work focused heavily on the Stochastic Block Model (SBM) and its variants, utilizing modularity maximization, maximum likelihood, and spectral clustering to recover block probabilities (Bickel and Chen, 2009; Abbe, 2017; Rohe et al., 2011). As the field moved toward continuous, nonparametric graphons, a number of estimators such as USVT (Chatterjee, 2015), sorting-based methods (Chan and Airoldi, 2014), and block-based estimators (Gao et al., 2015), achieve near-optimal or minimax rates under appropriate structural assumptions.
Among these, the neighborhood smoothing method introduced by Zhang et al. (2017) stands out. By adaptively aggregating structurally similar nodes based on empirical two-hop distances, neighborhood smoothing achieves minimax (or rate-optimal) MSE under their structural assumptions.
@@
This decoupling is the central theoretical lever of the paper. By strictly separating the data used for neighborhood selection from the data used for estimation, we restore exact conditional independence among the centered summands. This topological decoupling permits the application of classical empirical process tools to graphon neighborhood smoothing. Specifically, assuming the true graphon is dense and comprises a fixed number of well-separated, piecewise Lipschitz blocks, we leverage this restored independence to establish three main contributions:
@@
    Asymptotic Normality: We establish a uniform Berry-Esseen bound with explicit rates for the normalized variance component; normal-approximation intervals in this version use a heuristic variance estimate and incur an additional variance-estimation error term.
@@
    Honest Cross-Validation: The LOO structure naturally supports an unbiased cross-validation scheme for tuning the neighborhood size $h_{n}$. Here, honest cross-validation means that it does not reuse held-out edges in predictor construction, ensuring valid inference even after data-driven tuning.
@@
The remainder of the paper is organized as follows. Section 2 formalizes the graphon model, the Zhang-Levina-Zhu (ZLZ), Zhang et al. (2017) empirical distance, and our leave-one-out neighborhood construction. Section 3 establishes the foundational conditional independence, proving the variance-adaptive concentration bounds, Berry-Esseen normal approximations, and the validity of the resulting confidence intervals. Section 5 details the honest cross-validation scheme for tuning parameter selection. Section 6 reports comprehensive simulation results illustrating the finite-sample coverage and MSE performance, and Section 7 concludes with a discussion of extensions to directed, weighted, and degree-corrected settings.
@@ 2 Model and Leave-One-Out Neighborhood Smoothing
We observe a single undirected simple graph on $n$ vertices with adjacency matrix $A\in\{0,1\}^{n\times n}$, no self-loops ($A_{ii}=0$), and independent upper-triangular entries
@@
Assumption 2.2 places the model in the dense graphon regime, where edge probabilities are uniformly bounded away from zero and one. This excludes sparse scalings of the form $P_{ij}=\rho_{n}f(\xi_{i},\xi_{j})$ with $\rho_{n}\to 0$. Our Berry-Esseen and empirical Bernstein arguments rely critically on variance bounded away from zero. Extending the present theory to sparse or degree-corrected regimes would require variance renormalization and additional control of small-probability events, and is left for future work.
@@
Assumption 2.2 guarantees a uniformly nonvanishing conditional variance and is used to apply Berry-Esseen and variance-adaptive concentration bounds. It excludes extremely sparse regimes where $P_{ij}\to 0$ uniformly; extending our results to such regimes requires variance re-normalization and additional control of rare events, which we leave to future work.
@@
For concreteness we work in the canonical (dense) graphon regime $P_{ij}=f(\xi_{i},\xi_{j})$ with $f$ fixed as in Assumption 2.1. In this regime the expected degree satisfies $\mathbb{E}\deg(i)=\sum_{j=1}^{n}P_{ij}\asymp n$, so local neighborhoods contain many informative two-hop paths and several concentration steps simplify. All main results below are stated under Assumptions 2.1 and 2.2 together with this dense regime convention. Extensions to explicit sparsity scalings $P_{ij}=\rho_{n}f(\xi_{i},\xi_{j})$ are discussed briefly in Section 7.
@@ 2.1 LOO neighborhoods and estimator
Define the leave-one-out neighborhood of node $i$ by selecting the $h_{n}$ nearest neighbors according to $d^{(-j)}$. To ensure the neighborhood size is deterministically fixed, we resolve any distance ties at the boundary by prioritizing smaller node indices. Because the arbitrary integer labels assigned to nodes are independent of the latent positions $\xi$, this lexicographic tie-breaking is strictly independent of the underlying graphon geometry and introduces zero statistical bias. Formally, let $\pi$ be the permutation of $V\setminus\{i,j\}$ that sorts the nodes such that $d^{(-j)}(i,\pi(1))\leq d^{(-j)}(i,\pi(2))\leq\dots$, with $\pi(a)<\pi(b)$ if $d^{(-j)}(i,\pi(a))=d^{(-j)}(i,\pi(b))$. We define:
$\mathcal{N}_{i}^{(-j)}=\big\{\pi(1),\pi(2),\dots,\pi(h_{n})\big\}.$ (2.3)
Here, $h_{n}\asymp\sqrt{n\log n}$ is a fixed bandwidth sequence. Because $\sqrt{(n-1)\log(n-1)}\asymp\sqrt{n\log n}$, explicitly applying $h_{n}$ on the (n1)(n-1)-node reduced graph preserves the theoretical scaling of the classical ZLZ neighborhoods while strictly fixing the LOO neighborhood size to exactly $|\mathcal{N}_{i}^{(-j)}|=h_{n}$. Following the notations in Zhang et al. (2017), given $\mathcal{N}_{i}^{(-j)}$, define the one-sided LOO smoother and its symmetrized version by
@@
While the point estimator is practically symmetrized via $\widehat{P}_{ij}$ to respect the undirected nature of the graph, averaging the two distinct neighborhoods $\mathcal{N}_{i}^{(-j)}$ and $\mathcal{N}_{j}^{(-i)}$ reintroduces complex dependencies. However, because point estimation risk easily transfers via the inequality $(\widehat{P}_{ij}-P_{ij})^{2}\leq\frac{1}{2}(\tilde{P}_{ij}-P_{ij})^{2}+\frac{1}{2}(\tilde{P}_{ji}-P_{ij})^{2}$, the symmetric estimator inherits the optimal mean-squared error rate. Despite this advantage, valid entrywise uncertainty quantification requires exact conditional independence. Therefore, our inferential limit theorems and confidence intervals are developed and constructed strictly for the fundamental one-sided LOO predictor $\tilde{P}_{ij}$. Therefore, our inferential limit theorems and confidence intervals are developed and constructed strictly for the fundamental one-sided LOO predictor $\tilde{P}_{ij}$.
For any $i\neq j$, the entrywise estimation error of the one-sided LOO smoother can be additively decomposed into a stochastic fluctuation and a deterministic bias. We write
$\Delta_{ij}:=\tilde{P}_{ij}-P_{ij}=U_{ij}^{(-j)}+B_{ij}^{(-j)},$ (2.5)
where the variance component is defined as
$U_{ij}^{(-j)}:=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(A_{kj}-P_{kj}),$ (2.6)
and the deterministic bias component is defined as
Define the distance function $d:V\times V\to\mathbb{R}_{+}$ by
@@
This is the sup-norm discrepancy between the $i$th and $j$th rows of the normalized two-hop matrix $A^{2}/n$, measuring the similarity of their two-hop neighborhoods. Therefore the ZLZ neighborhood $\mathcal{N}_{i}$ consists of the $h_{n}$ nodes with the smallest ZLZ distance to $i$.
Following Zhang et al. (2017), let $q_{i}(h/n)$ denote the $(h/n)$-th empirical quantile of the distance set $\{d_{\mathrm{ZLZ}}(i,k):k\in V\setminus\{i\}\}$. The ZLZ neighborhood is defined as the set of nodes whose distance to $i$ falls below this quantile threshold:
$\mathcal{N}_{i}^{\mathrm{ZLZ}}:=\big\{k\in V\setminus\{i\}:d_{\mathrm{ZLZ}}(i,k)\leq q_{i}(h/n)\big\}.$ (2.9)
The ZLZ estimator is then defined by the empirical average over this neighborhood:
$\widetilde{P}^{\mathrm{ZLZ}}_{ij}:=\frac{1}{|\mathcal{N}_{i}^{\mathrm{ZLZ}}|}\sum_{k\in\mathcal{N}_{i}^{\mathrm{ZLZ}}}A_{kj}.$
Because $\mathcal{N}_{i}$ depends on all columns of $A$, including column $j$, the weights used to average $\{A_{kj}:k\in\mathcal{N}_{i}\}$ depend on the very summands being averaged. Hence the centered terms $A_{kj}-P_{kj}$ are not independent conditional on the neighborhoods, preventing the direct application of variance-adaptive concentration or Berry-Esseen theory.
@@
The symmetrized estimator $\widehat{P}_{ij}$ respects the symmetry of the simple graph model. All concentration and normal-approximation results are derived for the one-sided quantities $\tilde{P}_{ij}$; symmetrization does not affect the independence structure.
@@
Condition on the latent positions $(\xi_{1},\dots,\xi_{n})$ so that $P$ is deterministic with $P_{kl}=f(\xi_{k},\xi_{\ell})$. Under the model (2.1), the edges $\{A_{k\ell}:1\leq k<\ell\leq n\}$ are independent Bernoulli random variables with parameters $P_{k\ell}$. By construction, $A^{(-j)}$ is obtained from $A$ by removing the $j$th row and column. Hence $A^{(-j)}$ is a measurable function of $\{A_{k\ell}:k\neq j,\ell\neq j\}$, and so is $\mathcal{N}_{i}^{(-j)}$. Therefore, conditional on $P$, the random set $\mathcal{N}_{i}^{(-j)}$ is independent of the collection $\{A_{kj}:k\neq j\}$. For each $k\in\mathcal{N}_{i}^{(-j)}$, we have $A_{kj}\sim\mathrm{Bernoulli}(P_{kj})$ and $A_{kj}-P_{kj}\in[-1,1]$ with mean zero. Since the edges $\{A_{kj}:k\neq j\}$ are independent conditional on $P$, and $\mathcal{N}_{i}^{(-j)}$ is independent of these edges, it follows that the collection $\{A_{kj}-P_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ is independent, mean-zero, and bounded in [1,1][-1,1] conditional on $(\mathcal{N}_{i}^{(-j)},P)$. ∎
3 Main Results: Entrywise Uncertainty Quantification
With the conditional independence of the leave-one-out (LOO) summands established in Proposition 2.7, we are now equipped to construct valid, entrywise confidence intervals. We separate the estimation error into a stochastic variance component, governed by the conditionally independent edges, and a deterministic bias component, which captures the approximation error of the local neighborhood average.
3.1 Finite-Sample Confidence Intervals via Empirical Bernstein Bounds
To construct nonasymptotic confidence intervals for the variance component $U_{ij}^{(-j)}$, we require a concentration inequality that depends only on observable quantities. Classical Bernstein inequalities involve the true variance $V_{ij}$, which is unknown. Instead, we utilize the Empirical Bernstein bound of Maurer and Pontil (2009), which provides a deviation inequality that replaces the unknown variance with the sample variance, yielding a fully data-driven interval.
The unbiased sample variance of the observed edges in the LOO neighborhood is given by
$s_{ij}^{2}:=\frac{1}{h_{n}(h_{n}-1)}\sum_{\begin{subarray}{c}k,\ell\in\mathcal{N}_{i}^{(-j)}\\ k<\ell\end{subarray}}(A_{kj}-A_{\ell j})^{2},$ (3.1)
where $s_{ij}^{2}$ is the unbiased sample variance of the neighborhood observations. Because the observations $A_{kj}$ are binary indicators, this sample variance has the simplified exact form:
$s_{ij}^{2}:=\frac{1}{h_{n}(h_{n}-1)}\sum_{\begin{subarray}{c}k,\ell\in\mathcal{N}_{i}^{(-j)}\\ k<\ell\end{subarray}}(A_{kj}-A_{\ell j})^{2}=\frac{h_{n}}{h_{n}-1}\,\tilde{P}_{ij}(1-\tilde{P}_{ij}).$ (3.2)
where $\tilde{P}_{ij}=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj}$ is the empirical LOO mean.
Lemma 3.1 (Empirical Bernstein bound).
Assume $h_{n}\geq 2$. For any $\alpha\in(0,1)$, the stochastic fluctuation $U_{ij}^{(-j)}$, defined in (2.6), satisfies
$\mathbb{P}\!\left(\big|U_{ij}^{(-j)}\big|\leq\sqrt{\frac{2s_{ij}^{2}\log(4/\alpha)}{h_{n}}}+\frac{7\log(4/\alpha)}{3(h_{n}-1)}\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)\geq 1-\alpha.$ (3.3)